Blumen Textiles Corporation began April with a budget for 37,000 hours of production in the Weaving Department. The department has a full capacity of 49,000 hours under normal business conditions. The budgeted overhead at the planned volumes at the beginning of April was as follows:

Variable overhead $118,400
Fixed overhead 83,300
Total $201,700

The actual factory overhead was $204,100 for April. The actual fixed factory overhead was as budgeted. During April, the Weaving Department had standard hours at actual production volume of 38,000 hours. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number. Round your interim computations to the nearest cent, if required.

a. Determine the variable factory overhead controllable variance.
$   

b. Determine the fixed factory overhead volume variance.
$  
Answer #1

(1)

Budgeted rate of variable overhead = $118400/37000 = $3.2 per hour

Standard variable overhead for actual production = 38000 x $3.2 = $121600

Variable factory overhead controllable variance = Standard variable overhead - Actual variable overhead

= $121600 - ($204100 - $83300) = 800 Favorable

(2)

Predetermined fixed overhead rate = $83300/49000 = $1.70 per hour

Fixed overhead applied = standard hours x standard rate

= 38000 x $1.70 = $64600

Fixed overhead volume variance = Fixed overhead applied - Budgeted fixed overhead

= $64600 - $83300 = $18700 Unfavorable